Eva Report – January 4th

Crew 272 EVA Report 04-01-2023

EVA # 3

Author: Arly Black

Purpose of EVA: 1) Placement of the MicroFox 15 transmitter for the radio navigation system and drone scouting of Candor Chasma near Compass Rock (20 minutes) 2) Stratigraphy of Candor Chasma to begin interpreting depositional environments and gypsum sample collection for GreenHab experiments (2 hrs)

Start time: 1056

End time: 1340

Narrative: The EVA got off to a late start because of an issue with Adriana’s comms system which needed to be swapped out for another and troubleshooting of the GPS device. After a 25-minute delay (Madelyn wishes it to be noted that SHE was ready on time!) the intrepid explorers exited the airlock and began their journey… or tried to. Curiosity wouldn’t move – who knew that could happen when it was still plugged in?! Once that was resolved by the very complicated act of pulling out the charging cord, finally they were actually on their way down scenic Cow Dung Road. The snafus continued, however, as the crew soon found themselves lost on the ol Cow Dung. They missed their turn and had to backtrack 0.25 miles to Galileo Road. Their navigation was aided by an alien ship sighting near the destination of interest. About 0.25 miles down Galileo Road, Perseverance hit 60% charge, and while Curiosity was only at 66%, all four explorers parked their vehicles and continued the rest of the way on foot to a ledge that provided a good view of both Compass Rock and the Hab, where Kshitij began to test his transmitter system, known as the Boiler Transmission Station (BTS), as well as the transmitter left in the Hab. The transmitters were set to frequencies of 147 and 146 MHz, respectively. He was able to detect a loud tone from both transmitters using his Yagi-UDA antenna. The full plan was to drive to Candor Chasma and navigate back to BTS, but due to the poor SOC of the vehicles, this aspect of the EVA was unobtainable.

Simultaneously, the drone was tested by Mason. He flew it towards Candor Chasma but had connection issues while flying low to look for areas of interest, and the drone kept trying to return on its own. He will need to check the long-range beacon and try to set up the analog controller which should boost the range compared with using a phone. Mason was able, however, to get a cool orbital video of the EVA crew with the drone. At the ledge, Adriana helped Madelyn collect two containers of sand for her Mars regolith experiments. Although no gypsum was found, the mainly quartz sand and clay should be useful for her research.

Then everyone returned to the rovers, dropped of the drone and sandy samples, and continued along Galileo on foot until they reached the Compass Rock formation. There they found some good powdery sulfates to mix in with the sandy sample (useful as sulfates have been detected on Mars). Adriana collected some samples of petrified wood, sandstone, and some other unidentified minerals. She looked for paleo flow indicators but couldn’t find much evidence of ancient river systems. However, she saw evidence of modern fluvial processes, which could be useful for giving an indication of the ancient processes. The others had fun climbing the Rock! The crew decided not to go to Candor Chasma due to time and vehicle constraints. They returned to the rovers, got back on Cow Dung Road, and headed back to the Hab. On the way back, the EVA crew performed a rover speed test to see what led to increased battery drainage – turns out going slow is better for conservation of battery life. They additionally timed all walks throughout the expedition to make sure they could accurately budget how long they could stay in the field and marked the waypoints of BTS and Madelyn’s samples. They wish it to be noted that the entire EVA they thought and acted like true Martians!

Destination: Ledge near Galileo Road and Compass Rock

Coordinates (use UTM WGS 84): N519750, E4251500 (Boiler Ledge – self-named), N520000, E4252000 (Compass Rock)

Participants: Adriana Brown (Crew Geologist), Madelyn Whitaker (GreenHab Officer), Mason Kuhn (Crew Engineer), Kshitij Mall (Crew Commander)

Road(s) and routes per MDRS Map: Drove northeast on Cow Dung Road 0.25 miles past Galileo Road 1104. Backtracked to Galileo Road 1104 and followed road for 0.25 miles. Walked on foot along Galileo Road 1104 to N519750, E4251500. Returned to rovers, then walked back along Galileo Road 1104 to Compass Rock N520000, E4242000. Returned to vehicles, drove back to Cow Dung Road and then returned to Hab.

Mode of travel: Driving and walking

Operations Report – January 4th

Crew 272 Operations Report 4-1-2023

SOL: 3

Name of person filing report: Mason Kuhn

Non-nominal systems: Suit Charger #1, Hab toilet,

Notes on non-nominal systems: I swapped the battery of suit 1 to suit 2, and then charged suit 2 (battery 1) with charger 1. The voltage reading after about 2 hours was unchanged at 10.38 V. I then moved the battery of suit 1 back to suit 1. I took charger 11 (since nobody seems to be using the lightweight suit) and attached it to suit 1 (battery 1). After 2 hours the voltage reading of battery 1 was 13.37 V. From this, I believe the issue is with charger 1 and that there is no issue with suit 1.

No new issues to report with the Hab toilet.


Spirit rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Opportunity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: Yes

Hours: 218.9

Beginning charge: 100%

Ending charge: 59%

Currently charging: Yes

Perseverance rover used: Yes

Hours: 253.4

Beginning charge: 100%

Ending charge: 51%

Currently charging: Yes

General notes on rovers: Curiosity and Perseverance were used on the long EVA today. On our way out

Summary of Hab operations: Before the toilet tank was emptied, crewmates noticed bad-smelling liquid coming from below the toilet. I emptied the toilet, and then that liquid was cleaned and sanitized extra. This was a result of waiting a bit too long to empty the toilet.


Water (static tank): 331

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: All suits except suit 11 are plugged in. Charger 11 is being used for suit 1 because charger 1 was determined to be weak. All other suits are charging. Suits 7-10 were used on the EVA from 1030 to 1340 (returned 50 minutes early). All radios 1-7 are fully charged and have been removed from the charger.

Summary of GreenHab operations: Watered and monitored all GreenHab crops, no new plantings. Monitored experimental crops growing in paper substrate, set up new experiments using potting soil and mixture of substrate gathered on today’s EVA with bleached and unbleached seeds.

WATER USE: 7 gallons

Heater: On

Supplemental light: 4 hours

Harvest: 1g basil, 1g thyme, 1g oregano

Summary of ScienceDome operations:

Dual split: On

Summary of RAM operations: Retrieved a DMM and measuring tape around 0750 which have been in the lower deck of the Hab all day. I was checking suit charges and transmitter 9V battery charges throughout the day. The measuring tape has been used to track our water usage throughout the day (to find which operations use the most water) due to very high usage yesterday.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support:

  • Request for info on rover battery:
  • On today’s EVA, we noticed some strange battery data. On the way to our destination, both Perseverance and Curiosity were going around 80% speed on mode M (always in mode M), and the battery of each drained to 60-66% charge. We had to stop due to a rover reaching 60% charge and continue on foot. On our way back, both rovers were going around 60% speed on mode M (always in mode M), and the battery of each drained less than 10%. Going both directions, we slowed down significantly when the road terrain required slower speeds.

In short, is it normal for the rover battery to drain ~40% one way and ~10% on the way back? We were curious about if a slight speed decrease usually makes such a huge difference in battery usage.


Mason Kuhn

Crew Engineer

GreenHab Report – January 4th


Crew 272 GreenHab Report 04-01-2023

GreenHab Officer: Madelyn Whitaker

Environmental control: heater, fan

Average temperatures: 69 F, 20 C

Hours of supplemental light: 4 hours

Daily water usage for crops: 7 gallons

Daily water usage for research and/or other purposes: 0.5 gallons, taken from Hab sink

Water in Blue Tank : 281 gallons

Time(s) of watering for crops: 1600

Changes to crops: Harvested basil, thyme, oregano. No other changes.

Narrative: Watered and monitored all GreenHab crops, no new plantings. Monitored main experimental crops growing in paper substrate. Identified areas to add new crops next Sol.

Harvest: (include which crop and mass in grams): 1g basil, 1g thyme, 1g oregano

Support/supplies needed: N/A

Journalist Report – January 4th

Crew 272 Journalist Report 1-4-2023

Sol: 3

Author: Kenny Pritchard, Crew Journalist

Title: BTS

On the third sol of our stay at MDRS and the second of our simulation, we’re really breaking into our routine. Here at the Hab, everyone has their clear roles and responsibilities for the day, and everything gets done in a timely manner. It’s a simple lifestyle of chores, work, and relaxation that I think we could all get used to! I continued in my showcase of New Jersey cuisine this morning by following up last night’s pizzas with bagel/egg sandwiches for breakfast. Much of the crew elected to pass on the dehydrated eggs after their previous experience, but I’ll hold that they’re delicious with a healthy supplement of dehydrated cheese.

Adriana, Kshitij, Madelyn and Mason ventured out for a longer EVA today while the other three of us held down the Hab. This trip to the regions of Compass Rock and Candor Chasma was multi-faceted. The plan was for Mason to test scouting with his drone, Kshitij to set up a radio transmitter for his low-visibility navigation research, Madelyn to collect gypsum and regolith samples, and Adriana to fawn over seemingly ordinary rocks (and minerals!) as usual. According to our post-EVA debriefing session, it was a resoundingly partial success! Really though, the team did a good job adapting to the situation and accomplishing goals while the rest of us sat on our Martian behinds. Hiking around in EVA-attire, though fashionable, can be very tiring.

After nearly getting lost on the untamed roads around the Hab (seriously, Mars needs a transit administration), the EVA crew found their way and began their work. Kshitij was able to set up a beacon (the Boiler Transmission Station, or BTS) on a ledge overlooking Compass Rock. With the signal from this station and a nearly identical transmitter at the Hab, an astronaut could use a handheld ‘Yagi-Uda’ antenna to locate themselves without needing a visual. This might be important for a crew caught in one of those bothersome Martian dust storms I mentioned on Sol 1. Meanwhile, Mason was setting up a long-range beacon and piloting Garud the drone to scout ahead for the next destination of the EVA. Despite difficulties maintaining a solid connection at range, he was able to get some good views and an orbital video of the crew. Madelyn collected two containers of sand samples, which she will use as substrate for her plant growth experiments. The crew was ultimately unable to continue to Candor Chasma due to a combination of rover-battery and time concerns, but they pushed ahead to Compass Rock proper. There, Madelyn collected some good powdery sulfates, and Adriana some sandstone, petrified wood, and invertebrate fossils. Life on Mars! Somebody tell David Bowie.

Dinner was prefaced by a pungent odor emanating from the bathroom on the lower deck. Fear not, Mason suited up and fixed the problem, which required flushing out our septic system. Thanks Mason! It pays to have a good engineer out here. This time, he didn’t even die after making the repair. For dinner, Arly is cheffing up some cauliflower and sweet potato tacos, and we’ll celebrate our lack of sim-deaths this sol (so far…).

Sol Summary – January 4th


Crew 272 Sol Summary Report 04-01-2023

Sol: 3

Summary Title: Elusive Gypsum but finally BTS on Mars

Author’s name: Kshitij Mall

Mission Status: All nominal

Sol Activity Summary: After having a good sleep of 8 hours to digest the delicious Pizza prepared by Hemingway/Kenneth (recipe courtesy MDRS veteran Dr. Cesare Guariniello), the crew indulged in the daily routine of short warm up exercises lead by HSO Xerox (Megan). The crew then savored bagels of two flavors along with rehydrated eggs prepared by Hemingway and Xerox. Crew Engineer Scout (Mason) kept tabs on the water usage since the crew consumed more than nominal yesterday. After a briefing session at 0930, the EVA crew started preparing for the EVA. Due to minor anomalies in the GPS device and a minor hiccup with the radio comms, the EVA crew’s departure got slightly delayed. But since safety is of topmost priority, the EVA crew ensured it was completely ready to embark instead of hurrying up. On the side, I used the 3D printed toll made by Xerox to make a hole in the box to place the transmitter.

Thinking like a Martian was need of the hour and the EVA crew implemented it whenever a tough situation was presented. The EVA crew slightly got lost on the way after spotting (and possibly getting intimidated by) an alien ship comprising two aliens. After recovering the way, the EVA crew went ahead but the Perseverance rover reached 60% state of charge and the EVA crew chose to walk on foot to carry on the mission, The EVA crew the climbed on the top of a ledge that they described as the Boiler Ledge. At this ledge they placed a temporary transmission system comprising Microfox 15 transmitter antenna system of 147 MHz frequency. They named it the Boiler Transmission Station (BTS). I used a Yagi-Uda antenna to test the BTS as well a Microfox 15 transmitter placed in the Hab operating with 146 MHz frequency. While the BTS was being established, Scout started what his call sign exactly suggest: Scouting. For this he used a drone named Garud. The performance test of the drone was below nominal, so we counted it as a partial success. The EVA crew then embarked on Mission Gypsum on the approved EVA route but on foot to get some gypsum for the Greenhab Officer, Sprout (Madelyn). Crew Geologist Rocky (Adriana) served as the tour guide for Mars explaining about several Martian rock and river formations. Although gypsum eluded us throughout, other resources were collected to be used at the Greenhab. Rocky kept collecting rocks all along the way and the sampling bag was filled with plethora of rocks. Although the EVA crew had partial success in all the objectives, they felt nice that it was not a complete failure.

The EVA crew embarked on the journey back to the hab and this time tested the discharge rate of the battery on the rovers. We found that driving very slow without many acceleration changes indeed saved the charge on the rovers although more time was needed to cover the required distances. So, it will be a tradeoff. The EVA crew still made it back around 45 minutes before the requested arrival time. After coming back the whole crew filled in the cognitive performance test and the personal survey requested by the XO and Crew Scientist, Maple (Arly). By this time, it was also identified that there was no water leakage of any sorts in the hab, but it was indeed the crew using more water than nominal. The crew has taken a note of it and has taken all necessary actions to bring the water usage down to nominal. Scout and Rocky crashed to bed after an eventful EVA. But I had to forcibly wake up Scout to complete another mission: empty the toilet, which he successfully completed. Scout also completed vacuum cleaning the lower deck, while I am going to vacuum clean the upper deck. Hemingway and I practiced our Russian language skills for a small time, while venturing on some Spanish words on the side. Maple and Scout are on the dinner duty tonight and preparing sweet potato and cauliflower along with rice.

Look Ahead Plan: Continue the range test and scouting operations of Garud (our drone) and to see the BTS in action on Mars

Anomalies in work: None

Weather: Sunny, Clear, Winds Calm

Crew Physical Status: Good

EVA: One

Reports to be filed: Sol Summary, Operations Report, Journalist Report, EVA Report, EVA Request, Greenhab Report.

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site